Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

solve it

Expert replies
by maihuna » Fri Aug 14, 2009 11:26 am
Probability of solving specific problem independently by A and B are
1/2 and 1/3 respectively. If both try to solve the problem independently, find the probability that the problem is solved?

1/3
2/3
4/5
5/6
1
Charged up again to beat the beast :)
Join the discussion
Source: — Problem Solving |

by DanaJ » Fri Aug 14, 2009 11:46 am
First off, let's remember that the probability of something NOT happening is 1 - (the probability of that event happening).
Three cases in which the problem gets solved:

1. they both get it at the same time. P = 1/3 * 1/2 = 1/6

2. A gets it, but B does not. P = 1/2 * (1 - 1/3) = 1/2 * 2/3 = 1/3

3. A doesn't get it, but B does. P = (1 - 1/2)*1/3 = 1/2*1/3 = 1/6

Sum of P = 1/6 + 1/6 + 1/3 = 2/6 + 1/3 = 1/3 + 1/3 = 2/3.


Side note: another way of doing this is by noticing that the probability of the problem being solved is 1 - the probability of the event that the problem is not solved by either.

P' = (1 - 1/2)*(1 - 1/3) = 1/2 * 2/3 = 1/3

1 - P' = 1 - 1/3 = 2/3.
Join the discussion